OpenStudy (anonymous):

what is the answer to b -6b + 5=1 HELP PLEASE

OpenStudy (xhawaiiangirlx):

would u like me to show u how to solve it or just answer it straight out

OpenStudy (xhawaiiangirlx):

There are two answers for this decimal form and fraction form

pooja195 (pooja195):

Show your work and guide the user.

OpenStudy (xhawaiiangirlx):

refine the question

OpenStudy (xhawaiiangirlx):

5-5b=1

OpenStudy (xhawaiiangirlx):

Then subtract it by 5

OpenStudy (xhawaiiangirlx):

from both sides

OpenStudy (xhawaiiangirlx):

5-5b-5=1-5

OpenStudy (xhawaiiangirlx):

is what it should look like

OpenStudy (xhawaiiangirlx):

-5b=-4 is the outcome of that

OpenStudy (xhawaiiangirlx):

The divide both sides by -5

OpenStudy (xhawaiiangirlx):

-5b/-5 = -4/-5

OpenStudy (xhawaiiangirlx):

Then u get ur answer

OpenStudy (xhawaiiangirlx):

which is 4/5

OpenStudy (xhawaiiangirlx):

in decimal form it is 0.8

OpenStudy (xhawaiiangirlx):

have a nice day!